OpenStudy (arianna1453):

FAN +MEDAL Let g be a function that is defined for all x, x ≠ 2, such that g(3) = 4 and the derivative of g is g′(x) = x^2-16/x-2 with x ≠ 2. Find all values of x where the graph of g has a critical value. For each critical value, state whether the graph of g has a local maximum, local minimum, or neither. You must justify your answers with a complete sentence. On what intervals is the graph of g concave down? Justify your answer. Write an equation for the tangent line to the graph of g at the point where x = 3.

OpenStudy (arianna1453):

1. I got the critical values of x= 4 , -4 2. I got that since g''(x) >0 for x=4 , -4 that they are both local minimums. 3. I am stuck on finding the intervals of concave down.

OpenStudy (arianna1453):

@Astrophysics Can you help?

OpenStudy (arianna1453):

@mathmale ?

OpenStudy (arianna1453):

@ParthKohli

OpenStudy (mathmale):

"x ≠ 2" ?? Why the focus on x=2, if the function is defined for all x?

OpenStudy (mathmale):

This statment, "g′(x) = x^2-16/x-2 with x ≠ 2," is unusual. Why "with x=2?"

OpenStudy (arianna1453):

x cannot equal 2. Theres a vertical asymtote there.

OpenStudy (mathmale):

the important thing here is that you're given the derivative of function g(x). How do you go about using this derivative to determine the critical values of g(x)?

OpenStudy (mathmale):

OK. then write \[x \neq2\]

OpenStudy (arianna1453):

Ive already figure out the critical values. Its 4, -4

OpenStudy (mathmale):

This is the opposite of x=2, which was the reason I needed to ask for clarification.

OpenStudy (mathmale):

Please demonstrate that 4 and -4 are indeed critical values of g(x). How would you do that?

OpenStudy (arianna1453):

You set the derivative equal to 0. then solve for x. leading you to obviously get 4, -4

OpenStudy (mathmale):

So, if you substitute x=4 into g'(x), you should obtain what result?

OpenStudy (arianna1453):

You would get 0.

OpenStudy (mathmale):

Great. Let's move on.

OpenStudy (mathmale):

What do you need help with?

OpenStudy (arianna1453):

On what intervals is the graph of g concave down? Justify your answer.

OpenStudy (mathmale):

Starting with the given g '(x), find the second derivative, g ''(x).

OpenStudy (arianna1453):

\[g''(x)= \frac{ x ^{2}-4x+16 }{ (x-2)^{2} }\]

OpenStudy (mathmale):

I'm not going to check that. Given that you started with a rational function, your obtaining a rational function as the second derivative is approrpriate. Please set this second derivative = to 0 and solve for x. This is to set up intervals on which g(x) is concave up and on which it's concave down.

OpenStudy (mathmale):

Hints: If you set this g ''(x)=0 and want to solve for x, that's equivalent to setting the numerator = 0 and solving for x. Note that g ''(x) is not definied at x=2. Plot x=2 on a number line. If the roots of g ''(x) are real, also graph them on the same number line.

OpenStudy (arianna1453):

Thats the part I dont understand.

OpenStudy (arianna1453):

2+2i sqrt3 ? like i dont understand

OpenStudy (mathmale):

Set the numerator of this rational function = to 0. In other words,\[x^2-4x+16=0\] I see you've already done this and have solved for the roots. Your roots are "complex". No point in graphing them on the real x axis.

OpenStudy (mathmale):

the only relevant real number to graph on the x-axis is x=2. Choose an x value smaller than 2 and then another one greater than 2. Substitute these 2 numbers into the 2nd derivative, only for the purpose of determining the sign of the 2nd derivative at each value of x. Results?

OpenStudy (arianna1453):

I chose 3 and 1. Both god positive 13

OpenStudy (mathmale):

On the interval\[(-\infty,2),\] the second derivative is positive? negative? On the interval \[(2m \infty), the second der. is positive? \neg?\]

OpenStudy (mathmale):

So, your 2nd derivative is + on both sides of x=2. What does that tell you about the direction of concavity of each half of the graph?

OpenStudy (arianna1453):

Its concave up.

OpenStudy (arianna1453):

which means since the second derivative is always positive, it is never concave down, correct?

OpenStudy (mathmale):

yes. very good. However, I'd prefer you mention that x=2. the sec. deriv. is undefinited there, so I wouldn't want to say the curve of g(x) is concave up at x=2. Instead, use sets, as I did above.

OpenStudy (arianna1453):

Okay! Awesome. For the next step, Write an equation for the tangent line to the graph of g at the point where x = 3. Would it be: y-y1=m(x-x1) y-g(3) = g'(3)(x-3) y-4=-7(x-3)

OpenStudy (mathmale):

Yes, very good. You already have a formula for the slope of the t. l, and need only subst. x=3 into it to determine the slope of the tan. line at x=3. How do you propose to find the value of g(3)?

OpenStudy (arianna1453):

g(3) = 4 with the given information that the problem gave us. Then to find g'(3) I substituted 3 for x in g'(x) and got -7.

OpenStudy (mathmale):

I haven't attempted that. But you demonstrate considerable competence, so I'll go along with your result here.

OpenStudy (arianna1453):

so plugging it all into the equation, so far i got y-4 = -7 (x-3)

OpenStudy (mathmale):

Again, that strikes me as appropriate , although I haven't done the actual calculations myself. Nice going on your part!

OpenStudy (arianna1453):

Would I need to get it into y=mx+ form? or would that be the equation for the tangent line at x=3?

OpenStudy (arianna1453):

hello?

OpenStudy (mathmale):

that's be an appropriate form, the simplest you could choose. Sorry for the delay; I got sidetracked. Are you comfortable with the solution you've reached?

OpenStudy (arianna1453):

Yes, completely perfect, Thank you!